# Towards Understanding Variants of Invariant Risk Minimization through the Lens of Calibration

### v1

@@ Abstract
Machine learning models traditionally assume that training and test data are independently and identically distributed. However, in real-world applications, the test distribution often differs from training. This problem, known as out-of-distribution generalization, challenges conventional models. Invariant Risk Minimization (IRM) emerges as a solution, aiming to identify features invariant across different environments to enhance out-of-distribution robustness. However, IRM’s complexity, particularly its bi-level optimization, has led to the development of various approximate methods. Our study investigates these approximate IRM techniques, employing the Expected Calibration Error (ECE) as a key metric. ECE, which measures the reliability of model prediction, serves as an indicator of whether models effectively capture environment-invariant features. Through a comparative analysis of datasets with distributional shifts, we observe that Information Bottleneck-based IRM, which condenses representational information, achieves a balance in improving ECE while preserving accuracy relatively. This finding is pivotal, as it demonstrates a feasible path to maintaining robustness without compromising accuracy. Nonetheless, our experiments also caution against over-regularization, which can diminish accuracy. This underscores the necessity for a systematic approach in evaluating out-of-distribution generalization metrics, one that beyond mere accuracy to address the nuanced interplay between accuracy and calibration.
@@ 1 Introduction
In the evolving landscape of machine learning, the importance of out-of-distribution (OOD) generalization (Arjovsky, 2020) and calibration (Guo et al., 2017) cannot be understated, especially in real-world applications and critical scenarios. Traditional machine learning approaches, grounded in the assumption that data are independently and identically distributed (IID), often struggle to cope with OOD scenarios. Additionally, a notable trend in contemporary machine learning models is their tendency towards overconfidence, which undermines the reliability of their confidence estimations. This issue is primarily recognized as a calibration challenge (Guo et al., 2017).
Responding to these limitations, there has been a surge in research focusing on methodologies like Invariant Risk Minimization (IRM) (Arjovsky et al., 2020). IRM is designed to identify features that remain constant across different environments, thereby facilitating more robust generalization in the face of environmental variations. However, the computational demands of IRM have led to the exploration of more feasible, approximate methods. Despite the development of several IRM variants, their inability to consistently surpass the performance of finely-tuned Empirical Risk Minimization (ERM) models has been observed (Gulrajani & Lopez-Paz, 2020)(Zhang et al., 2023). This performance gap raises critical questions about the inherent deficiencies of these IRM variants. As we showed in Figure 1, even by leveraging IRM variants, there is a trade-off between in-distribution accuracy and OOD accuracy. It implies that models trained by IRM variants failed to obtain invariant features since they sacrifice in-distribution accuracy to achieve better OOD accuracy.
From another perspective of IRM, it can also be conceptualized as a specialized form of multi-domain calibration (Wald et al., 2021). Our study intersects with these findings, examining the extent to which IRM variants mimic the original model in terms of calibration. We posit that effective implementation of IRM should lead to reduced loss and a zero expected calibration error (ECE) (Wald et al., 2021). Through comparative experimentation, we assess various IRM variants in environments marked by distributional shifts, such as diversity and correlation shifts (Ye et al., 2021). Our evaluation focuses not only on accuracy but also on ECE. The results highlight notable disparities in ECE among different methodologies, even when OOD accuracy levels are similar. This suggests that a sole reliance on OOD accuracy for method development might be inadequate. Our observations include a range of calibration behaviors across methods, from overconfidence to underestimation.
@@
  • We investigate the reasons behind the gap between IRM formulation and its variants through the lens of ECE.
  • We demonstrate that Information Bottleneck-based IRM (IB-IRM) effectively lowers ECE, aligning it more closely with the original objectives of IRM (Figure 3, Table 2).
@@
Utilizing ECE as a metric, our study sheds light on the gap between the original formulation of IRM and its variants. We recognize a notable trade-off between ECE and accuracy, highlighting the complexities and necessary compromises in striving for both robust generalization and precise calibration.
ID IRMv1 OOD IRMv1
ERM IRMv1 BIRM IB-IRM IRM Game PAIR
Figure 1: A graph observed in CMNIST showing the trade-off between in-distribution (ID) and OOD accuracy in approximation methods of IRM. Figure (a) visualizes the accuracy of the IRMv1 model ID and OOD, plotted against the training steps on the horizontal axis. The two solid lines are approximately symmetric around an accuracy of 50%. Figure (b) visualizes the ID accuracy (horizontal axis) against the OOD accuracy (vertical axis) for typical approximation methods of IRM. There is a clear trend that as the accuracy improves OOD, the ID accuracy decreases.
@@ 2.1 OOD Generalization
In traditional deep neural network scenarios, it is assumed that the learning and test environments are IID. The most commonly used learning algorithm in this assumption has been ERM(Vapnik, 1991), which has been successful in various scenarios. Specifically, ERM aims to minimize the sum of losses of the model $f:\mathcal{X}\to\mathcal{Y}$ in the training environment $E_{train}$.
@@
However, in the real world, the assumption of IID does not always hold, and distribution shifts between training and testing environments can occur (Quionero-Candela et al., 2009). When the model relies on features that are only effective in the training environment, known as shortcut features, its performance can decrease in the testing environment (Geirhos et al., 2020). This issue is commonly known as OOD generalization and has become an urgent problem in recent years. Therefore, it is necessary to train robust models that are not dependent on specific environments and can perform consistently across all environments. A robust model is defined as follows.
@@ 2.2 Calibration
The alignment of a model’s predictive confidence with its actual accuracy is extremely important in practical scenarios such as medical image diagnosis and autonomous driving. For example, if there are 100 instances where a model predicts with 80% confidence, then 80 of those predictions should be correct. However, in recent deep neural networks, issues such as increased model capacity have led to problems with the miscalibration and many methods have been proposed to address this calibration issue (Guo et al., 2017). The correct calibration in all environments is defined for any predictive probability $\alpha$ of $f$ as follows:
@@
where the predictive probabilities ranging from 0 to 1 are divided into $M$ bins, and for each bin, $Acc$ represents the accuracy of predictions within the bin, while $Conf$ represents the average predictive probability (confidence level) within the bin.
@@ 2.3 OOD Generalization and Calibration
It has been pointed out that OOD generalization and calibration are interrelated. Notably, (3) can be a sufficient condition for (2), theoretically demonstrating that calibration across all environments leads to OOD generalization.
Additionally, it has been empirically demonstrated that there is a correlation between a model’s accuracy in OOD (OOD) scenarios and its calibration performance (Ovadia et al., 2019).
In this paper, we utilize this fact and hypothesize that models with high calibration performance are likely to achieve better OOD generalization. We evaluate the approximation methods of Invariant Risk Minimization (IRM), introduced in 4, based on the model’s calibration performance to assess their OOD generalization capabilities.
@@ 3.1 Domain Invariant Representation
In the context of OOD scenarios, there is an application for Representation Learning, specifically aimed at extracting environment-invariant features from raw data and reducing environment-specific spurious features to achieve more robust predictions. For example, in a task of classifying animals from images, if the model learns to associate parts of an animal with extraneous elements, it may lead to high performance within the training domain but fail to make robust predictions when the background distribution shifts in the test domain. However, if the model discards spurious features (e.g., background) from the raw data, it can consistently make robust predictions across all environments. One particularly well-known method that applies representation learning to out-of-distribution generalization scenarios is Domain-Adversarial Neural Networks (DANN) (Ganin et al., 2016). This approach uses adversarial training to adjust the learning process so that it is impossible to discern from the extracted features which environment the input comes from, thereby aiming to extract invariant features.
@@ 3.2 Invariant Learning
The formulation in (5) represents a bi-level optimization problem, which is challenging to solve exactly. As a result, various implementable approximation methods have been proposed. However, despite numerous contributions to the development of approximation methods for IRM, a method that generally surpasses well-tuned ERM in terms of out-of-distribution generalization accuracy has not yet been proposed. Therefore, it is essential to understand why these approximation methods have not succeeded, and what approaches might be more suitable for addressing the approximation problem.
@@ 4 Related Works
Table 1 presents an overview of IRM and its approximation methods. We discuss the variants of IRM in the following section.
@@
IRM $\Phi$ $\omega$
IRMv1 $\omega\cdot\Phi$ 1.0
IB-IRM $\omega\cdot\Phi$ 1.0
PAIR $\omega\cdot\Phi$ 1.0 (Pareto Optimality)
@@
BIRM $\omega\cdot\Phi$ 1.0
@@ 4.2 Information Bottleneck based IRM(IB-IRM)
While IRMv1 has been successful with certain OOD generalization datasets such as CMNIST, it has been noted that if not enough invariant features are included in each environment in the training data, robust prediction is lost. Therefore, IB-IRM(Ahuja et al., 2022) aims to improve IRMv1 by compressing the entropy of $\Phi$ using the information bottleneck method, thereby preventing excessive reliance on features that adversely affect the invariant prediction. The variance of the features extracted from each data by $\Phi$ is added as a regularization term.
$\lambda\parallel\nabla_{\omega^{\prime}\mid\omega^{\prime}=1.0}R^{e}(\omega^{\prime}\cdot\Phi^{\prime})\parallel^{2}+\ \gamma Var(\Phi)$ (7)
@@ 4.3 Pareto IRM(PAIR)
In general, there is a trade-off between ERM and OOD generalization, and PAIR(Chen et al., 2023) focuses on the need to properly manage that trade-off when generalizing a model out of distribution. To achieve more robust models, PAIR is designed to find their Pareto optimal solutions in terms of multi-objective optimization. In fact, it is implemented as a multi-objective optimal problem for ERM, IRMv1 penalties, and vREx(Krueger et al., 2021) penalties.
@@ 5 Experiments
Although the approximation methods in the previous chapter have achieved some success in terms of OOD accuracy on OOD datasets (Arjovsky et al., 2020)(Ahuja et al., 2022)(Chen et al., 2023)(Ahuja et al., 2020)(Lin et al., 2022), it is not known how invariant features they actually acquire and how robust they are to unknown environments. Therefore, we evaluated them by comparing their calibration performance with ECE.
@@ 5.2.1 Dataset
CMNIST is a dataset based on MNIST, containing 70,000 examples of dimension (1, 28, 28). It is adapted for a binary classification task where digits less than 5 are labeled as class 0, and those 5 or above as class 1. Spurious correlations are added by assigning colors (red or green) to the digits, with the proportion of the two colors varying in each class depending on the environment. Denoting the proportion of green in class 0 as environment $e$, then the training environments are set as $e=[0.1,0.2]$, and the test environment as $e=[0.9]$.
@@ 5.2.2 Results
ERM IRMv1 BIRM IB-IRM IRM Game PAIR 100% PREDS 50% PREDS
Figure 2: Comparison of OOD accuracy (horizontal axis) and OOD ECE (vertical axis) in CMNIST. The red solid line represents the theoretical values when all model’s predictive probabilities are 100%, and the blue solid line represents the case when all predictive probabilities are 50%. IRMv1, IB-IRM, and BIRM, which achieve relatively high inference accuracy, are distributed near the blue line. This indicates that these models mitigate reliance on non-invariant features by making predictions more ambiguous.
@@
$\lambda$ in IRMv1
Figure 3: Comparison of the model’s OOD ECE in CMNIST, with various values of $\lambda$ in the formulation of IRMv1 6. It was observed that increasing $\lambda$, thereby increasing the regularization penalty in IRMv1, results in a lower ECE for the model in OOD environment.
Fig. 2 presents a scatter plot showing the relationship between OOD accuracy (horizontal axis) and ECE (vertical axis) for various methods, each trained with multiple hyperparameters. The red solid line represents the theoretical value when all model’s predictive probabilities are 100%, indicating very high confidence, while the blue solid line represents the case when all predictive probabilities are 50%, indicating very low and ambiguous confidence. The methods with relatively high accuracy (IRMv1(green), IB-IRM(light blue), BIRM(red)) align with the blue solid line, showing that they can overcome excessive fitting to non-invariant features by making predictions more ambiguous. On the other hand, ERM(blue) and IRM Game(purple) tend to be distributed near the red solid line, indicating predictions with relatively high confidence. PAIR(yellow) is distributed between the two solid lines, demonstrating characteristics that are intermediate between the two aforementioned groups.
Figure 3 visualizes how the penalty introduced in IRMv1 for the purpose of out-of-distribution generalization actually affects the ECE. It specifically compares different values of $\lambda$ in IRMv1’s formulation 6, demonstrating that as the impact of IRMv1’s penalty increases, the ECE decreases, suggesting better calibration. Additionally, the effects of the two regularization penalties in IB-IRM are similarly shown in the Appendix, Figures 6 and 7. In each case, comparisons were made by varying the magnitude of the regularization coefficient of the focused penalty while keeping the other penalty sufficiently influential. These results, like those in Figure 3, suggest that each penalty contributes to the improvement of the model’s ECE.

### v2

@@ Abstract
Machine learning models traditionally assume that training and test data are independently and identically distributed. However, in real-world applications, the test distribution often differs from training. This problem, known as out-of-distribution (OOD) generalization, challenges conventional models. Invariant Risk Minimization (IRM) emerges as a solution that aims to identify invariant features across different environments to enhance OOD robustness. However, IRM’s complexity, particularly its bi-level optimization, has led to the development of various approximate methods. Our study investigates these approximate IRM techniques, employing the Expected Calibration Error (ECE) as a key metric to measure the extent to which the model has acquired invariant features. ECE, which measures the reliability of model prediction, serves as an indicator of whether models effectively capture environment-invariant features. Through a comparative analysis of datasets with distributional shifts, we observe that Information Bottleneck-based IRM, which condenses representational information, achieves a balance in improving ECE while preserving accuracy relatively. This finding is pivotal, demonstrating a feasible path to maintaining robustness without compromising accuracy. Nonetheless, our experiments also caution against over-regularization, which can diminish accuracy. This underscores the necessity for a systematic approach in evaluating OOD generalization metrics, which goes beyond mere accuracy to address the nuanced interplay between accuracy and calibration. Our code is available at https://github.com/katoro8989/IRM_Variants_Calibration
@@ 1 Introduction
In the evolving landscape of machine learning, the importance of out-of-distribution (OOD) generalization (Vapnik, 1991) and calibration (Guo et al., 2017a) cannot be understated, especially in real-world applications and critical scenarios. Traditional machine learning approaches, grounded in the assumption that data are independently and identically distributed (IID), often struggle to cope with OOD scenarios. Additionally, a notable trend in contemporary machine learning models is their tendency towards overconfidence, which undermines the reliability of their confidence estimations. This issue is primarily recognized as a calibration challenge. Responding to these limitations, there has been a surge in research focusing on methodologies like Invariant Risk Minimization (IRM) (Arjovsky et al., 2020). IRM is designed to identify constant features across different environments, facilitating more robust generalization in the face of environmental variations. However, the computational demands of IRM have led to the exploration of more feasible, approximate methods. Despite the development of several IRM variants, their inability to consistently surpass the performance of finely-tuned Empirical Risk Minimization (ERM) models has been observed (Gulrajani & Lopez-Paz, 2020)(Zhang et al., 2023a). This performance gap raises critical questions about the inherent deficiencies of these IRM variants. As we showed in Figure 1, even by leveraging IRM variants, there is a trade-off between in-distribution (ID) and OOD accuracy. It implies that models trained by IRM variants failed to obtain invariant features since they sacrifice ID accuracy to achieve better OOD accuracy. From another perspective of IRM, it can also be conceptualized as a specialized form of multi-domain calibration (Wald et al., 2021a). Our study intersects with these findings, examining how IRM variants mimic the original model in calibration. We posit that effective implementation of IRM should lead to reduced loss and a zero expected calibration error (ECE) (Wald et al., 2021a). Through comparative experimentation, we assess various IRM variants in environments marked by distributional shifts, such as diversity and correlation shifts (Ye et al., 2021). Our evaluation focuses not only on accuracy but also on ECE. The results highlight notable disparities in ECE among different methodologies, even when OOD accuracy levels are similar. This suggests that a sole reliance on OOD accuracy for method development might be inadequate. Our observations include a range of calibration behaviors across methods, from overconfidence to underestimation.
ID IRMv1 OOD IRMv1 ERM IRMv1
BIRM IB-IRM IRM Game PAIR
Figure 1: A graph observed in CMNIST showing the trade-off between ID and OOD accuracy in approximation methods of IRM. Figure (a) visualizes the accuracy of the IRMv1 model ID and OOD, plotted against the training steps on the horizontal axis. The two solid lines are approximately symmetric around an accuracy of 50%. Figure (b) visualizes the ID accuracy (horizontal axis) against the OOD accuracy (vertical axis) for typical approximation methods of IRM. There is a clear trend that as the accuracy improves OOD, the ID accuracy decreases.
@@
  • We used ECE to analyze discrepancies between IRM formulation and its variants.
  • We observed that Information Bottleneck-based IRM (IB-IRM) effectively lowers ECE, aligning it more closely with the original objectives of IRM (Figure 3, Table 2).
@@
Our study, using ECE as a metric, illuminates the gap between the original formulation of IRM and its variants. We recognize a notable trade-off between ECE and accuracy, highlighting the complexities and necessary compromises in striving for robust generalization and precise calibration.
@@ 2.1 OOD Generalization
In traditional deep neural network scenarios, it is assumed that the learning and test environments are IID. The most commonly used learning algorithm in this assumption has been ERM(Vapnik, 1991), which has been successful in various scenarios. Specifically, ERM aims to minimize the sum of the risk of the model $f:\mathcal{X}\to\mathcal{Y}$ in the set of training environments $E_{train}$.
@@
where $R^{e}(f)$ is represented as $\mathbb{E}_{X^{e},Y^{e}}[\ell(f(X^{e}),Y^{e})]$, which is the risk in a environment $e$.
However, in the real world, the assumption of IID does not always hold, and distribution shifts between training and testing environments can occur(Quinonero-Candela et al., 2008). When the model relies on features that are only effective in the training environment, known as shortcut features, its performance can decrease in the testing environment(Geirhos et al., 2020). This issue is commonly known as OOD generalization and has become an urgent problem recently. Therefore, it is necessary to train robust models that are not dependent on specific environments and can perform consistently across all environments. A robust model is defined as follows.
@@ 2.2 Calibration
The alignment of a model’s predictive confidence with its actual accuracy is extremely important in practical scenarios such as medical image diagnosis and autonomous driving. For example, if there are 100 instances where a model predicts with 80% confidence, then 80 of those predictions should be correct. However, in recent deep neural networks, issues such as increased model capacity have led to problems with miscalibration, and many methods have been proposed to address this calibration issue (Guo et al., 2017a). The correct calibration in all environments is defined for any predictive probability $\alpha$ of $f$ as follows:
@@
where the predictive probabilities ranging from 0 to 1 are divided into $M$ parts, $Acc(B_{m})$ is the prediction accuracy within the bin $B_{m}$ and $Conf(B_{m})$ is the average of the prediction probability $\hat{p}_{b}$ (confidence level) for each data in the bin $B_{m}$.
@@ 2.3 OOD Generalization and Calibration
It has been pointed out that OOD generalization and calibration are interrelated. Notably, (3) can be a sufficient condition for (2), theoretically demonstrating that calibration across all environments leads to OOD generalization. Wald et al. (2021b) capitalize on this observation and propose CLOvE, a method that calibrates models across multiple environments, thereby enabling generalization to unseen data distributions.
Additionally, it has been empirically demonstrated that there is a correlation between a model’s accuracy in OOD scenarios and its calibration performance. It has been demonstrated that there exists a correlation between a model’s accuracy and its ECE under dataset shift scenarios. Furthermore, techniques such as temperature scaling(Guo et al., 2017b), originally proposed for calibration purposes, have been shown to be effective for OOD generalization.(Ovadia et al., 2019; Naganuma & Hataya, 2023).
In this paper, we utilize this fact and hypothesize that models with high calibration performance will likely achieve better OOD generalization. We evaluate the approximation methods of Invariant Risk Minimization (IRM), introduced in 4, based on the model’s calibration performance to assess their OOD generalization capabilities.
@@ 3.1 Domain Invariant Representation
In the context of OOD scenarios, there is an application for Representation Learning, specifically aimed at extracting environment-invariant features from raw data and reducing environment-specific spurious features to achieve more robust predictions. For example, in a task of classifying animals from images, if the model learns to associate parts of an animal with extraneous elements, it may lead to high performance within the training domain but fail to make robust predictions when the background distribution shifts in the test domain. However, if the model discards spurious features (e.g., background) from the raw data, it can consistently make robust predictions across all environments. One particularly well-known method that applies representation learning to OOD generalization scenarios is Domain-Adversarial Neural Networks (DANN) (Ganin et al., 2016). This approach uses adversarial training to adjust the learning process so that it is impossible to discern from the extracted features which environment the input comes from, thereby aiming to extract invariant features.
@@ 3.2 Invariant Learning
The formulation in (5) represents a bi-level optimization problem, which is challenging to solve exactly. As a result, various implementable approximation methods have been proposed. Many variants add various constraints to the IRM objective and approximate the bi-level optimization problem as a single-level problem to make it more tractable (Arjovsky et al., 2020; Ahuja et al., 2022; Chen et al., 2022; Ahuja et al., 2020; Lin et al., 2022). Additionally, Zhang et al. (2023b) propose an improved variant of IRM Game, another IRM variant, that retains the bi-level optimization framework during training. However, despite numerous contributions to the development of approximation methods for IRM, a method that generally surpasses well-tuned ERM in terms of OOD generalization accuracy has not yet been proposed. Therefore, it is essential to understand why these approximation methods have not succeeded, and what approaches might be more suitable for addressing the approximation problem.
While the original IRM formulation poses implementation challenges, a alternative method has also been proposed to address this issue. IRM aims to learn invariant features by keeping $\mathbb{E}[Y^{e}|f(X^{e})]$ constant across environments. However, Huh & Baidya (2023) introduced a complementary notion of invariance called MRI, which seeks to conserve the label-conditioned feature expectation $\mathbb{E}[f(X^{e})|Y^{e}]$ across environments, in contrast to the original IRM formulation.
@@ 4 Related Works
IRM
$\Phi$
$\omega$
IRMv1
$\omega\cdot\Phi$
1.0
IB-IRM
$\omega\cdot\Phi$
1.0
PAIR
$\omega\cdot\Phi$
1.0
IRM Game
$\Phi$
$\omega_{ens}$
@@
BIRM
$\omega\cdot\Phi$
1.0
@@ 4.2 Information Bottleneck based IRM(IB-IRM)
While IRMv1 has been successful with certain OOD generalization datasets such as CMNIST, it has been noted that if not enough invariant features are included in each environment in the training data, robust prediction is lost. Therefore, IB-IRM(Ahuja et al., 2022) aims to improve IRMv1 by compressing the entropy of the feature extractor $\Phi^{\prime}$ using the information bottleneck method, thereby preventing excessive reliance on features that adversely affect the invariant prediction. The variance of the features extracted from each data by $\Phi^{\prime}$ is added as a regularization term.
$\lambda\parallel\nabla_{\omega^{\prime}\mid\omega^{\prime}=1.0}R^{e}(\omega^{\prime}\cdot\Phi^{\prime})\parallel^{2}+\ \gamma Var(\Phi^{\prime})$ (7)
@@ 4.3 Pareto IRM(PAIR)
In general, there is a trade-off between ERM and OOD generalization, and PAIR(Chen et al., 2022) focuses on the need to properly manage that trade-off when generalizing a model out of distribution. To achieve more robust models, PAIR is designed to find their Pareto optimal solutions in terms of multi-objective optimization. In fact, it is implemented as a multi-objective optimal problem for ERM, IRMv1 penalties, and VREx(Krueger et al., 2021) penalties.
@@ 4.5 Bayesian IRM(BIRM)
ERM IRMv1 BIRM IB-IRM
IRM Game PAIR 100% Confidence 50% Confidence
Figure 2: Comparison of OOD accuracy (horizontal axis) and OOD ECE (vertical axis) in CMNIST. The red solid line represents the theoretical values when all the model’s predictive probabilities are 100%, and the blue solid line represents the case when all predictive probabilities are 50%. IRMv1, IB-IRM, and BIRM, which achieve relatively high inference accuracy, are distributed near the blue line. This indicates that these models mitigate reliance on non-invariant features by making predictions more ambiguous.
@@ 5 Experiments
Although the approximation methods in the previous chapter have achieved some success in terms of OOD accuracy on OOD datasets (Arjovsky et al., 2020; Ahuja et al., 2022; Chen et al., 2022; Ahuja et al., 2020; Lin et al., 2022), it is not known how invariant features they actually acquire and how robust they are to unknown environments. Therefore, we evaluated them by comparing their calibration performance with ECE.
@@ 5.2.1 Dataset
CMNIST is a dataset based on MNIST, containing 70,000 dimension examples (1, 28, 28). It is adapted for a binary classification task where digits less than 5 are labeled as class 0, and those 5 or above as class 1. Spurious correlations are added by assigning colors (red or green) to the digits, with the proportion of the two colors varying in each class depending on the environment. Denoting the proportion of green in class 0 as environment $e$, then the training environments are set as $e=[10\%,20\%]$, and the test environment as $e=[90\%]$.
@@ 5.2.2 Results
Figure 3: Comparison of the model’s OOD ECE in CMNIST, with various values of $\lambda$ in the formulation of IRMv1 6. It was observed that increasing $\lambda$, thereby increasing the regularization penalty in IRMv1, results in a lower ECE for the model in the OOD environment.
Fig. 2 presents a scatter plot showing the relationship between OOD accuracy and ECE for multiple IRM variants, each trained with multiple hyperparameters. The red solid line represents the theoretical value when all model’s predictive probabilities are 100%, indicating very high confidence, while the blue solid line represents the case when all predictive probabilities are 50%, indicating very low and ambiguous confidence. The methods with relatively high accuracy (IRMv1(green), IB-IRM(light blue), BIRM(red)) align with the blue solid line, showing that they can overcome excessive fitting to non-invariant features by making predictions more ambiguous. On the other hand, ERM(blue) and IRM Game(purple) tend to be distributed near the red solid line, indicating predictions with relatively high confidence. PAIR(yellow) is distributed between the two solid lines, demonstrating characteristics that are intermediate between the two aforementioned groups.
Figure 3 visualizes how the penalty introduced in IRMv1 for the purpose of OOD generalization actually affects the ECE. It specifically compares different values of $\lambda$ in IRMv1’s formulation 6, demonstrating that as the impact of IRMv1’s penalty increases, the ECE decreases, suggesting better calibration.